Can some one please help me to solve these radical expression and help me to decide if they have a solution. If the radical has no solution can you please explain why? Thanks in advance for your help.

1. Solve. Sqrt x=-3 If the equation does not have a solution, explain why.

2. Solve. Sqrt x+5= Sqrt 2x+8. If the equation does not have a solution, explain why.

3. Solve. x-3= Sqrt 17-4x. If the equation does not have a solution, explain why.

Answer by lwsshak3(11628)

Can some one please help me to solve these radical expression and help me to decide if they have a solution. If the radical has no solution can you please explain why? Thanks in advance for your help.
1. Solve. Sqrt x=-3 If the equation does not have a solution, explain why.
√x=-3
square both sides
x=9
..
2. Solve. Sqrt x+5= Sqrt 2x+8. If the equation does not have a solution, explain why.
√(x+5)=√(2x+8)
x+5=2x+8
2x-x=5-8
x=-3
..
3. Solve. x-3= Sqrt 17-4x. If the equation does not have a solution, explain why.
(x-3)=√(17-4x)
square both sides
x^2-6x+9=17-4x
x^2-2x-8=0
factor
(x-4)(x+2)=0
x=4
or
x=-2